OpenStudy (anonymous):

Help plz!!! f(x)=x^2+x-4 and g(x)=x+3 find h(x)=(f o g)(x)

OpenStudy (anonymous):

@Aylin

OpenStudy (anonymous):

h(x)=(f o g)(x)=f[g(x)] f[g(x)]=[g(x)]^2+g(x)-4 f[g(x)]=(x+3)^2+x+3-4 f[g(x)]=x^2+6x+9+x+3-4 h(x)=f[g(x)]=x^2+7x+8 do you understand??